OpenStudy (anonymous):

In how many ways can 9 instructors be assigned to six sections of a course in mathematics?

OpenStudy (anonymous):

9 x 6 = 54 ways

OpenStudy (anonymous):

\[\left(\begin{matrix}9 \\ 6\end{matrix}\right)\] = 84

OpenStudy (anonymous):

I don't know if this is right I think that i have to go 9*8*7*6*5*4*3*2*1=362880 It's probability so i have no idea

OpenStudy (amistre64):

9.8.7.6.5.4

OpenStudy (anonymous):

\[_{9}P_{6}=9!/(9-6)!=9*8*7*6*5*4\] Basically this is a permutation. Pretending that the classrooms are boxes that we are holding still, we have 9 professors, so the question is how many different ways can we order the professors (equivalent to sticking them in each of the classrooms) taking them six at a time?